OpenStudy (happykiddo):

what are the real or imaginary solutions of the polynomial equation::::: x^3-8=0

OpenStudy (luigi0210):

When it's cubed there are no imaginary solutions x^3=8 x=2

OpenStudy (rajee_sam):

This is difference of cubes form. (a³ - b³) = (a -b) ( a² + ab + b² ) x³ - 8 = x³ - 2³ Now use the above formula to write your factors.

OpenStudy (rajee_sam):

Then find your roots